john can finish his homework on 10 minutes. a fellow student would finish it up at a different rate. both students working together can do the homework in 8 minutes. how long does it take the second student to complete the homework if working alone.
Answer by jorel1380(3719)

1/10+1/x=1/8
1/x=1/8-1/10
1/x=5/40-4/40
1/x=1/40
x=40
It takes the second student 40 minutes working alone to finish the homework..